Joe has 37 coins (all nickels, dimes, and quarters) worth $5.50. He has 4 more quarters than nickels. How many dimes does Joe have?

Answer by stanbon(75887)

Joe has 37 coins (all nickels, dimes, and quarters) worth $5.50. He has 4 more quarters than nickels. How many dimes does Joe have?
================
Let the number of nickels be "x"; Value of those coins is 5x cents.
The number of quarters is "x+4" ; Value of those is 25(x+4)=25x+100 cents
And number of dimes is "37-x-(x+4)"="33-2x"; Value is 10(33-2x)=330-20x cents
EQUATION:
value + value + value = 550 cents
5x+25x+100+330-20x=550
10x=120
x=12 (number of nickels)
x+4=16 (number of quarters)
37-28=9(number of dimes)
Cheers,
Stan H.